# Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes

### v1

@@ Abstract
Open optical cavity resonators coupled to quantum emitters (such as atoms, molecules, and quantum dots) are a fundamental building block for many modern quantum technologies such as nanolasers [1, 2, 3], quantum sensors [4, 5], quantum computers [6, 7, 8, 9], and quantum networks [10, 11, 12, 13]. For a theoretical treatment of such systems, knowledge of the full electromagnetic Green’s function is of great importance and a requirement to many quantum optics formalisms. For example, it is used to calculate the Purcell enhancement of optical cavities [14, 15, 16, 17], in the quantization of the electromagnetic field [18, 19, 20, 21, 16, 22], and in the derivation of coupling elements between photons and quantum emitters [23, 24]. Therefore, an efficient scheme for obtaining a good approximation of the full Green’s function is central to many theoretical treatments of light-matter coupling in such systems.
One approach to open cavity systems is through quasinormal modes (QNMs), which are solutions to the source-free Helmholtz equation under open boundary conditions [25, 26, 27, 15]. The QNMs have complex eigenfrequencies with a negative imaginary part, so that the temporal decay is inherent to the mode, making the QNMs the natural modes of open cavity systems with material losses.
@@
In this Letter, we introduce a powerful and accurate Dyson equation approach to obtain the multi-cavity scattered Green’s function from only single-cavity QNMs, within a few-mode approximation. We use regularized QNM fields outside the individual resonators to characterize the intercavity scattering, making the approach applicable to systems with large spatial separations between the cavities, where retardation effects are non-negligible. Multi-cavity scattering naturally decomposes into products of two-cavity scattering processes within the QNM expansion, avoiding nested integrals. We calculate the coupling between two dipole emitters located in the gaps of metal dimers serving as QNM cavities, and compare the QNM expansion with the full numerical Green’s function to excellent agreement.
@@
where $\epsilon_{\rm back}(\mathbf{r},\omega)$ is the permittivity of the background medium, $V_{i}(\mathbf{r},\omega)=\chi_{\mathcal{V}_{i}}(\mathbf{r})\big[\epsilon(\mathbf{r},\omega)-\epsilon_{\rm back}(\mathbf{r},\omega)\big]$ is the perturbation of the permittivity due to the presence of the cavity, with $\mathcal{V}_{i}$ the cavity volume, and the function $\chi_{\mathcal{V}_{i}}(\mathbf{r})$ is unity for $\mathbf{r}\in\mathcal{V}_{i}$ and zero elsewhere.
@@
under appropriate boundary conditions. For cavities in vacuum, as considered in this paper, this is the Silver-Müller radiation condition [35, 36].
@@
Now, $\mathbf{G}^{(n)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ can be obtained from the Green’s function $\mathbf{G}^{(n-1)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ for n1n-1 cavities via the scattering equation [see App. A (in end matter) for the derivation]
@@
where $\mathcal{V}_{n}$ is the volume of the $n$-th cavity. For $n=1$ and using $\mathbf{G}^{(0)}\equiv\mathbf{G}_{\rm back}$, we recover the Dyson equation for the single-cavity case from Ref. 28.
For any fixed $n=k$, Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) is recursive, since $\mathbf{G}^{(k)}$ appears also on the right-hand side (RHS) of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes). However, this $\mathbf{G}^{(k)}$ on the RHS only depends on positions within the same cavity  [^1] For any fixed $n=k$, either (i) $\mathbf{r}^{\prime}\in\mathcal{V}_{k}$, or (ii) since $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime},\omega)=[\mathbf{G}(\mathbf{r}^{\prime},\mathbf{r},\omega)]^{T}$ alternatively $\mathbf{r}\in\mathcal{V}_{k}$ will recover case (i) by using the transpose of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), or (iii) if neither $\mathbf{r}^{\prime}$ nor $\mathbf{r}$ are in $\mathcal{V}_{k}$ inserting the transpose of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) on the RHS of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) again recovers case (i).. Consequently, we can terminate the Dyson equation at low order for any $k$ by replacing $\mathbf{G}^{(k)}$ at the RHS with the single-cavity QNM Green’s function ($1\leq i\leq k$)
@@
i.e., we assume that the Green’s function for positions inside the cavity can be approximated by the single-cavity Green’s function of that cavity [cf. Eq. (4)]. This assumption holds well for many applications, such as 3D cavities in homogeneous media, and other conditions for termination or perturbative treatments of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) may be viable for cases where the assumption breaks down.
As a consequence, the full Green’s function for $N$ cavities, $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime},\omega)=\mathbf{G}^{(N)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ is obtained iteratively from a set of $N$ scattering equations (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) for n=1,...,Nn=1,...\,,N (cf. Fig. 1).
As we show for three coupled cavities in App. C, the ordering of the cavities 1,...,N1,...\,,N should be performed such that the Dyson equation (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) terminates [i.e., the case from Eq. (8) is obtained on the RHS] at low order for each step in the iteration.
This formalism yields the scattered Green’s function for an arbitrary number of cavities within a finite number of iterations. In addition, the QNM expansion factorizes the Green’s function [cf. Eq. (4)], so that multi-cavity scattering processes naturally decompose into products of two-cavity scattering processes (see App. C for an example of three-cavity scattering), thus avoiding numerically demanding nested integrals.
For a practical and efficient numerical calculation, the volume integral representation from Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) can be converted into surface integrals, yielding (omitting the $\omega$-dependence in the Green’s functions for a brief notation, cf. App. A):
@@
where, $j=1,2$, and $\hbar\omega_{\rm p}=8.2934\,{\rm eV}$ is the plasma frequency, while $\hbar\gamma_{\rm p1}=0.0928\,{\rm eV}$ and $\hbar\gamma_{\rm p2}=0.3\hbar\gamma_{\rm p1}$. The QNM of dimer $1$ has the eigenfrequency $\hbar\tilde{\omega}_{1}=(1.6904-i0.0652)\,{\rm eV}$, corresponding to a wavelength of $\lambda_{1}\approx 733.46\,{\rm nm}$. For the QNM of dimer $2$, we obtain $\hbar\tilde{\omega}_{2}=(1.6482-i0.0388)\,{\rm eV}$ and $\lambda_{2}\approx 752.24\,{\rm nm}$.
@@
as an indicator to compare the QNM expansion from Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) to the full Green’s function [46].
From Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), we obtain (see App. B)
@@
Note that $B_{21}(\omega)=B_{12}(\omega)$ holds (see App. B), preserving $\mathbf{G}(\mathbf{r}_{a},\mathbf{r}_{b},\omega)=[\mathbf{G}(\mathbf{r}_{b},\mathbf{r}_{a},\omega)]^{T}$.
@@
with $N_{21}=(2.0694-i0.1357)\cdot 10^{-7}\,{\rm nm}^{-2}$ [cf. Eq. (End Matter) in App. C].
In Fig. 3, we show the imaginary and real parts of $g_{ba}(\omega)$, calculated using the full numerical Green’s function and the QNM approximation from Eq. (12) together with Eq. (14), and observe excellent agreement. We stress that we do not use any fitting parameters in the calculations. The frequency-dependent phase $\mathrm{e}^{i\omega R_{12}/c}$ from Eq. (14) is crucial for cases with significant separation. We show in Fig. 3 the case where the exponential is included in the pole approximation ($G^{\rm QNM}_{\rm pole}$),
@@
Figure 3: Normalized coupling $g_{ba}(\omega)$ [cf. Eq. (11)] between the two dipoles in Fig. 2 for $R_{12}=2020\,{\rm nm}$. The QNM expansion from Eq. (12) together with Eq. (14) ($G^{\rm QNM}$) yields excellent agreement with the full numerical Green’s function calculation ($G^{\rm full}$). Using Eq. (15) instead of Eq. (14) ($G^{\rm QNM}_{\rm pole}$) does not fully include retardation effects. The single-cavity QNM expansion ($G_{1}$) [Eq. (16)] does not match the full numerical results, confirming the coupling of the QNMs. The coupling is normalized to $g^{\rm back}_{ba}(\omega_{1})$ (i.e., without the dimers).
For comparison, we also show in Fig. 3 a single-cavity expansion of the Green’s function [cf. Eq. (18)]
@@
In Fig. 4, we show the coupling $g_{ba}(\omega)$ for a shorter dimer separation of $R_{21}=760\,{\rm nm}$, which is close to the QNM resonance wavelengths. The QNM expansion from Eq. (12) together with Eq. (14) [where $N_{21}=(5.3494+i0.2984)\cdot 10^{-7}\,{\rm nm}$] again shows excellent agreement with full numerical calculations. Including the phase in the pole approximation [Eq. (15)] yields better agreement with the full solution than for $R_{21}=2020\,{\rm nm}$, since retardation effects are weaker for shorter separations. The single-cavity expansion [Eq. (16)] again fails to capture the full coupling, both quantitatively and qualitatively.
Figure 4: Same as in Fig. 3, but with $R_{12}=760\,{\rm nm}$. Equation (12) together with Eq. (14) ($G^{\rm QNM}$) agrees excellently with the full numerical solution ($G^{\rm full}$). Equation (15) ($G^{\rm QNM}_{\rm pole}$) yields better agreement than for larger separations, while the single-cavity QNM expansion ($G_{1}$) from Eq. (16) fails to capture the shape and magnitude of the coupling.
@@
Comparing to rigorous numerical solutions of the full 3D Maxwell equations to the QNM expansion, we demonstrated the quantitative accuracy of the approach for two coupled metal dimers serving as QNM cavities at two different distances between the dimers.

### v2

@@ Abstract
Introduction.—Open optical cavity resonators coupled to quantum emitters (such as atoms, molecules, and quantum dots) are a fundamental building block for many modern quantum technologies such as nanolasers [3, 39, 43], quantum sensors [37, 26], quantum computers [33, 6, 5, 7], and quantum networks [9, 34, 10, 36]. For a theoretical treatment of such systems, knowledge of the full electromagnetic Green’s function is of great importance and a requirement to many quantum optics formalisms. For example, it is used to calculate the Purcell enhancement of optical cavities [49, 44, 12, 40], in the quantization of the electromagnetic field [19, 11, 46, 35, 12, 16], and in the derivation of coupling elements between photons and quantum emitters [15, 28]. Therefore, an efficient scheme for obtaining a good approximation of the full Green’s function is central to many theoretical treatments of light-matter coupling in such systems.
One approach to open cavity systems is through quasinormal modes (QNMs), which are solutions to the source-free Helmholtz equation under open boundary conditions [27, 30, 25, 44]. The QNMs have complex eigenfrequencies with a negative imaginary part, so that the temporal decay is inherent to the mode, making the QNMs the natural modes of open cavity systems with material losses. Quasinormal mode theories have been used to study a variety of setups in recent works [44, 50, 51, 38, 18, 29], and play a crucial role in rigorous open cavity quantum optics [19, 12, 16].
@@
In this Letter, we introduce a powerful and accurate Dyson equation approach to obtain the multi-cavity scattered Green’s function for systems with large spatial separations and non-negligible retardation from only single-cavity QNMs, within a few-mode approximation. We use regularized QNM fields outside the individual resonators to characterize the intercavity scattering, so that retardation effects are included. Multi-cavity scattering naturally decomposes into products of two-cavity scattering processes within the QNM expansion, avoiding nested integrals. This approach allows for a few-mode expansion of the multi-cavity Green’s function, and thus the implementation in quantum optics applications. We calculate the coupling between two dipole emitters located in the gaps of metal dimers serving as QNM cavities, and compare the QNM expansion with the full numerical Green’s function to excellent agreement.
@@
where $\epsilon_{\rm back}(\mathbf{r},\omega)$ is the permittivity of the background medium, $V_{i}(\mathbf{r},\omega)=\chi_{\mathcal{V}_{i}}(\mathbf{r})\big[\epsilon(\mathbf{r},\omega)-\epsilon_{\rm back}(\mathbf{r},\omega)\big]$ is the perturbation of the permittivity due to the presence of the cavity, with $\mathcal{V}_{i}$ the cavity volume, and the function $\chi_{\mathcal{V}_{i}}(\mathbf{r})$ is unity for $\mathbf{r}\in\mathcal{V}_{i}$ and zero elsewhere. We make no assumption about the background permittivity $\epsilon_{\rm back}$ except those necessary to define QNMs, i.e., that there is a discontinuity between the permittivity of the cavity medium and $\epsilon_{\rm back}$, and that there is no backscattering of waves from the far field [8].
@@
under appropriate boundary conditions, which is sufficient for defining the Green’s function mathematically. For cavities in vacuum, as considered in this paper, this is the Silver-Müller radiation condition [31, 45].
@@
Now, $\mathbf{G}^{(n)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ can be obtained from the Green’s function $\mathbf{G}^{(n-1)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ for n1n-1 cavities via the scattering equation [see S2 A in Ref. 1 for the derivation]
@@
where $\mathcal{V}_{n}$ is the volume of the $n$-th cavity. For $n=1$ and using $\mathbf{G}^{(0)}\equiv\mathbf{G}_{\rm back}$, we recover the Dyson equation for the single-cavity case from Ref. 17.
For any fixed $n=k$, Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) is recursive, since $\mathbf{G}^{(k)}$ appears also on the right-hand side (RHS) of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes). While a perturbative treatment of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) is generally possible (cf. Sec. S2 C in [1]), this $\mathbf{G}^{(k)}$ on the RHS only depends on positions within the same cavity  [^1] For any fixed $n=k$, either (i) $\mathbf{r}^{\prime}\in\mathcal{V}_{k}$, or (ii) if $\mathbf{r}\in\mathcal{V}_{k}$, case (i) is recovered by using the transpose of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), or (iii) if neither $\mathbf{r}^{\prime}$ nor $\mathbf{r}$ are in $\mathcal{V}_{k}$ inserting the transpose of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) on the RHS of Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) again recovers case (i).. Consequently, we can terminate the Dyson equation at low order for any $k$ by replacing $\mathbf{G}^{(k)}$ at the RHS with the single-cavity QNM Green’s function ($1\leq i\leq k$)
@@
i.e., we assume that the Green’s function for positions inside the cavity can be approximated by the single-cavity Green’s function of that cavity [cf. Eq. (4)]. This assumption holds well for many applications, such as 3D cavities in homogeneous media (where the influence of other cavities scales as $1/R$) with separations on the order of the QNM wavelength or more (see Fig. 4 as well as Figs. S2 and S3 in Ref. 1). For cases where the assumption breaks down, the perturbative approach (see S2 C in [1]) or the established CQT are employed.
As a consequence, the full Green’s function for $N$ cavities, $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime},\omega)=\mathbf{G}^{(N)}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ is obtained iteratively from a set of $N$ scattering equations (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) for n=1,,Nn=1,...\,,N (cf. Fig. 1).
As we show for three coupled cavities in S2 D of Ref. 1 the ordering of the cavities 1,,N1,...\,,N should be performed such that the Dyson equation (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) terminates [i.e., the case from Eq. (8) is obtained on the RHS] at low order for each step in the iteration.
This formalism yields the scattered Green’s function for an arbitrary number of cavities within a finite number of iterations. In addition, the QNM expansion factorizes the Green’s function [cf. Eq. (4)], so that multi-cavity scattering processes naturally decompose into products of two-cavity scattering processes (see S2 D in Ref. 1 for an example of three-cavity scattering), thus avoiding numerically demanding nested integrals.
For a practical and efficient numerical calculation, the volume integral representation from Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) can be converted into surface integrals. Towards, this, we note that from Eq. (6), it follows that $\Delta\epsilon(\mathbf{s},\omega)\big|_{\mathbf{s}\in\mathcal{V}_{n}}=\epsilon^{(n)}(\mathbf{s},\omega)-\epsilon^{(n-1)}(\mathbf{s},\omega)$ in Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes). Thus, we employ the Helmholtz equation (2) with $\epsilon^{(n)}$ for $\mathbf{G}^{(n)}$ and with $\epsilon^{(n-1)}$ for $\mathbf{G}^{(n-1)}$ in Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), to find via the second Green’s theorem (omitting the $\omega$-dependence in the Green’s functions for a brief notation, see Sec. S2 B in Ref. 1 for details):
@@
where, $j=1,2$, and $\hbar\omega_{\rm p}=8.2934\,{\rm eV}$, $\hbar\gamma_{\rm p1}=0.0928\,{\rm eV}$ are the plasma frequency and damping rate for gold, while dimer 2 has a reduced damping rate of $\hbar\gamma_{\rm p2}=0.3\hbar\gamma_{\rm p1}$ to obtain a higher-quality plasmon mode. The QNM of dimer $1$ has the eigenfrequency $\hbar\tilde{\omega}_{1}=(1.6904-i0.0652)\,{\rm eV}$, corresponding to a wavelength of $\lambda_{1}\approx 733.46\,{\rm nm}$. For the QNM of dimer $2$, we obtain $\hbar\tilde{\omega}_{2}=(1.6482-i0.0388)\,{\rm eV}$ and $\lambda_{2}\approx 752.24\,{\rm nm}$.
@@
as an indicator to compare the QNM expansion from Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes) to the full Green’s function, which is obtained numerically from the scattered electric field of a point dipole source (see S1 in [1]).
To obtain $\mathbf{G}(\mathbf{r}_{b},\mathbf{r}_{a},\omega)$ from Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), we use $\mathcal{S}_{n}\to\mathcal{S}_{1}$ so that $\mathbf{s},\mathbf{r}_{a}\in\mathcal{V}_{1}$, allowing us to apply the termination condition from Eq. (8). We also employ the single-cavity QNM expansion [15]
$\mathbf{G}^{(n-1)}(\mathbf{s},\mathbf{r}_{b},\omega)$ $\to\mathbf{G}_{2}(\mathbf{s},\mathbf{r}_{b},\omega)\big|_{\mathbf{s}\notin\mathcal{V}_{2},\mathbf{r}_{b}\in\mathcal{V}_{2}}$
$=A_{2}(\omega)\tilde{\mathbf{F}}_{2}(\mathbf{s},\omega)\tilde{\mathbf{f}}_{2}(\mathbf{r}_{b}).$ (12)
Thus, we arrive at
@@
Note that $B_{21}(\omega)=B_{12}(\omega)$ holds, preserving $\mathbf{G}(\mathbf{r}_{a},\mathbf{r}_{b},\omega)=[\mathbf{G}(\mathbf{r}_{b},\mathbf{r}_{a},\omega)]^{T}$ for reciprocal media.
@@
with
$N_{21}$ $=\oint_{\mathcal{S}_{1}}\mathrm{d}A_{s}\Big\{\Big[\nabla_{s}\times\tilde{\mathbf{F}}_{2}^{\prime}(\mathbf{s},\omega_{2})\Big]\cdot\Big[\hat{\mathbf{n}}_{s}\times\tilde{\mathbf{f}}_{1}(\mathbf{s})\Big]$
$\qquad\qquad-\Big[\hat{\mathbf{n}}_{s}\times\tilde{\mathbf{F}}_{2}^{\prime}(\mathbf{s},\omega_{2})\Big]\cdot\Big[\nabla_{s}\times\tilde{\mathbf{f}}_{1}(\mathbf{s})\Big]\Big\}.$ (16)
In Fig. 3, we show the imaginary and real parts of $g_{ba}(\omega)$, calculated using the full numerical Green’s function and the QNM approximation from Eq. (13) together with Eq. (15) $N_{21}=(2.0773-i0.0657)\cdot 10^{-7}\,{\rm nm}^{-2}$], and observe excellent agreement. We stress that we do not use any fitting parameters in the calculations. The frequency-dependent phase $\mathrm{e}^{i\omega R_{12}/c}$ from Eq. (15) is crucial for cases with significant separation. We show in Fig. 3 the case where the exponential is included in the pole approximation ($G^{\rm QNM}_{\rm pole}$),
@@
Figure 3: Normalized coupling $g_{ba}(\omega)$ [cf. Eq. (11)] between the two dipoles in Fig. 2 for $R_{12}=2020\,{\rm nm}\approx 2.75\lambda_{1}$. The QNM expansion from Eq. (13) together with Eq. (15) ($G^{\rm QNM}$) yields excellent agreement with the full numerical Green’s function calculation ($G^{\rm full}$). Using Eq. (17) instead of Eq. (15) ($G^{\rm QNM}_{\rm pole}$) does not fully include retardation effects. The single-cavity QNM expansion ($G_{1}$) [Eq. (18)] does not match the full numerical results, confirming the coupling of the QNMs. The coupling is normalized to $g^{\rm back}_{ba}(\omega_{1})$ (i.e., without the dimers).
For comparison, we also show in Fig. 3 a single-cavity expansion of the Green’s function [cf. Eq. (Green’s function expansion for multiple coupled optical resonators with finite retardation using quasinormal modes), $1\leftrightarrow 2$]
@@
In Fig. 4, we show the coupling $g_{ba}(\omega)$ for a shorter dimer separation of $R_{21}=760\,{\rm nm}$, which is close to the QNM resonance wavelengths. The QNM expansion from Eq. (13) together with Eq. (15) [where $N_{21}=(5.3797+i0.4523)\cdot 10^{-7}\,{\rm nm}$] again shows excellent agreement with full numerical calculations. Including the phase in the pole approximation [Eq. (17)] yields better agreement with the full solution than for $R_{21}=2020\,{\rm nm}$, since retardation effects are weaker for shorter separations. The single-cavity expansion [Eq. (18)] again fails to capture the full coupling, both quantitatively and qualitatively.
Figure 4: Same as in Fig. 3, but with $R_{12}=760\,{\rm nm}\approx 1.04\lambda_{1}$. Equation (13) together with Eq. (15) ($G^{\rm QNM}$) agrees excellently with the full numerical solution ($G^{\rm full}$). Equation (17) ($G^{\rm QNM}_{\rm pole}$) yields better agreement than for larger separations, while the single-cavity QNM expansion ($G_{1}$) from Eq. (18) fails to capture the shape and magnitude of the coupling.
@@
Comparing rigorous numerical solutions of the full 3D Maxwell equations to the QNM expansion, we demonstrated the quantitative accuracy of the approach for two coupled metal dimers serving as QNM cavities at two different distances between the dimers.